Use the Shell Method to compute the volume of the solids obtained by rotating the region enclosed by the graphs of the functions y=x^2 , y=8−x^2 and to the right of x=0.5 about the y-axis.


1
Expert's answer
2021-10-06T00:30:59-0400

"y = x^2, y = 8 − x^2,and \ x = 1/2" about the y-axis.


Here is how set up the integral:


"2 \pi \int [-2,2] (1/2 - x)[(8-x^2)- (x^2)]dx"


Since x=1/2 is greater than a (a= -2), I subtracted x from 1/2.


Also, since the graph of"y= 8- x^2" is above the graph of"y=x^2," Isubtracted x^2 from 8-x^2.